OpenStudy (anonymous):

ΔTOY has coordinates T (-3, 4), O (-4, 1), and Y (-2, 3). A translation maps point T to T' (-1, 1). Find the coordinates of O' and Y' under this translation.

OpenStudy (anonymous):

... the difference in coordinates between points T and T' determines the shifting or translation of axes... the resulting difference applies on other two points O and Y to O' and Y' respectively...

OpenStudy (anonymous):

shift of axes (x and y) = -3 - (-1) = -2 units going to left, and 4 - 1 = 3 units going down... therefore the coordinates of O' will be -4 + 2 = -2 for x, 1 - 3 = -2 for y or O' (-2,-2)...

OpenStudy (anonymous):

i will leave the Y' for you as an exercise...